Consider the following information for James Inc. for the Year 2014:

2014
Assets
Cash 55,600
Accounts Receivable 140,400
Inventory 185,400
Land 90,000
PP&E 225,000
Accumulated Depreciation (117,000)
Total Assets 579,400
Liabilities
Accounts Payable 151,200
Accrued Expenses 36,000
Notes Payable 52,000
Total Liabilities 239,200
Stockholders' Equity
Common Stock 18,000
Additional Paid-in Capital 115,200
Beginning Retained Earnings 145,800
Net Income 97,200
Dividends 36,000
Ending Retained Earnings 207,000
Total Stockholders' Equity 340,200
Total Liabilities and Equity 579,400

1. Calculate James' Debt-to-Equity Ratio

     

2. Calculate James' Debt Ratio
Answer #1

1;debt to equity ratio=Total liabilities/Total equity

=(239200/340200)=0.7031(Approx).

2.Debt ratio=Total liabilities/Total assets

=(239200/579400)=0.4128(Approx).